OpenStudy (anonymous):

Use the quadratic formula to solve the equation. –4x2 – 3x + 2 = 0

OpenStudy (anonymous):

4x2 – 3x + 2 = 0 or 4x2 + 3x - 2 = 0 (multiplying - on both sides) or 16x2 +12x -8 = 0 (multiplying 4 on both sides) or (4x)2 +2*4x*3/2+(3/2)2-(3/2)2-8=0 or (4x+3/2)2=(3/2)2+8=9/4+8=41/4 or 4x+3=sqrt(41/4)=(sqrt41)/2 or 4x=-3+(sqrt41)/2 or x=-3/4+(sqrt41)/8 ans

OpenStudy (anonymous):

\[x =\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a } \]

OpenStudy (anonymous):

here a=-4 , b = -3 and c = 2 you will get two values of x that will be the required solution.

OpenStudy (anonymous):

okay, thanks a bunch!